OpenStudy (anonymous):

Factor completely and check mentally 10a^2+20a

OpenStudy (anonymous):

this is alot of steps

OpenStudy (anonymous):

Well? All I know to say is take out a...

OpenStudy (anonymous):

Simplifying 10a2 + 20a = 0 Reorder the terms: 20a + 10a2 = 0 Solving 20a + 10a2 = 0 Solving for variable 'a'. Factor out the Greatest Common Factor (GCF), '10a'. 10a(2 + a) = 0 Ignore the factor 10. Subproblem 1 Set the factor 'a' equal to zero and attempt to solve: Simplifying a = 0 Solving a = 0 Move all terms containing a to the left, all other terms to the right. Simplifying a = 0 Subproblem 2 Set the factor '(2 + a)' equal to zero and attempt to solve: Simplifying 2 + a = 0 Solving 2 + a = 0 Move all terms containing a to the left, all other terms to the right. Add '-2' to each side of the equation. 2 + -2 + a = 0 + -2 Combine like terms: 2 + -2 = 0 0 + a = 0 + -2 a = 0 + -2 Combine like terms: 0 + -2 = -2 a = -2 Simplifying a = -2 Solution a = {0, -2}

OpenStudy (anonymous):

(0-2) is your answer

OpenStudy (anonymous):

hope i helped adios! aie aie aie!

OpenStudy (muzzack):

Final result : \(\huge\frak\ 10a • (a + 2)\)

OpenStudy (anonymous):

Omg.. Idk which one to go with.